OpenStudy (anonymous):

A sample of N2 in a flask at 25'C exerts a pressure of 525 torr. When 0.1 g of 02 is added to this flask, the pressure rises to 760 torr. The temperature constant and there is no reaction between N2 and O2. How many grams of N2 in the flask?

OpenStudy (mos1635):

N2: P1*V=n1*R*T N2+O2: P2*V=(n1+n2)*R*T divide P1/P2=n1/n1+n2 (n2=m2/MM2) solve to n1 convert to gr

OpenStudy (anonymous):

Thanks a lot @mos1635 :)